OpenStudy (anonymous):

uh how to do this x^3.x^2=x^6 the solution pls.......

OpenStudy (anonymous):

some typo here

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

\[x^3\times x^2=x^6\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the x^3 X^2 the exponent will be multiplied how bout the 2 x's ??

OpenStudy (anonymous):

ok then \[x^5=x^6\] \[x^6-x^5=0\] \[x^5(x-1)=0\] \[x^5=0\implies x = 0\] or \[x-1=0\implies x = 1\]

OpenStudy (anonymous):

what should be the real answer??